1. A jet plane has a takeoff speed of ? = 75 m/s and can move along the runway at an average acceleration of 1.3 m/s2. If the length of the runway is 2.5 km, will the plane be able to use this runway safely? Defend your answer.
2. A person takes a trip, driving with a constant speed of 89.5 km/h, except for a 22.0 min rest stop. If the person’s average speed is 77.8 km/h, (a) how much time is spent of the trip and (b) how far does the person travel?
3. A ball is thrown directly downward with an initial speed of 8.00 m/s, from a height of 30.0 m. After what time interval does it strike the ground?
Answer #1

Solution) (1) v = 75 m/s

a = 1.3 m/s^2

L = 2.5 km

v^2 = u^2 + 2aS

(75)^2 = 0 + 2(1.3)(S)

S = (75^2)/(2.6)

S = 2,163.5 m

S = 2.163 Km

Since S < L , so the jet can safely use this runway

* kindly post remaining questions in next post